OpenStudy (anonymous):

B^-4=1/81 b=? @yummydum

OpenStudy (anonymous):

\[b^-4=1/81\]

OpenStudy (anonymous):

Is it possible for there to be more than one answer?

OpenStudy (anonymous):

Nope @ Najia2000

OpenStudy (anonymous):

Oh, so do you have answer choices?

OpenStudy (anonymous):

No

OpenStudy (anonymous):

anything to the negative exponent means it's in the opposite position, so if something with a negative variable is in the numerator, it goes as a positive exponent in the denominator and vice versa so we have:\[b^{-4}={1\over81}\]meaning:\[{1\over b^4}={1\over81}\]so essentially:\[b^4=81\]now we have to figure out the fourth root of 81... which is 3:\[3\times3\times3\times3=81\]\[3^4=81\]\[81=81\checkmark\]so:\[{1\over3^4}={1\over81}\] however, do note that since the exponent is even there is a second answer, \(-3\)\[b=\pm3\]

OpenStudy (freckles):

hint: \[81=3^{4}\]

OpenStudy (anonymous):

so b= positive 3

OpenStudy (anonymous):

or \[+3\]

OpenStudy (freckles):

@yummydum gave you all the answers

OpenStudy (freckles):

well all the real ones anyways :p